OpenStudy (xmoses1):

The x-intercept of the line whose equation is 2x - 3y = 6 is 3 6 -2

OpenStudy (xmoses1):

I think that it is -2

jimthompson5910 (jim_thompson5910):

Plug in y = 0 and solve for x 2x - 3y = 6 2x - 3*0 = 6 2x - 0 = 6 2x = 6 x = 6/2 x = 3 So the x-intercept is 3, which is the point (3,0) Note: what you found (which is -2) is actually the y-intercept.

OpenStudy (xmoses1):

Oh so i plugged in the wrong variable

OpenStudy (xmoses1):

I have one more but i want to try and solve and you can tell me if i am correct

OpenStudy (xmoses1):

The y-intercept of the line whose equation is 2x + 5y = 8 is 8 4 8/5

jimthompson5910 (jim_thompson5910):

ok go for it

OpenStudy (xmoses1):

8/5??

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (xmoses1):

cool! thanks @jim_thompson5910 you helped me out a lot!

jimthompson5910 (jim_thompson5910):

you're welcome